TABLES

TABLE I. Average of the square of the magnetization for different lattice sizes L using the KBD algorithm with no Metropolis sweeps, for four different initial zero temperature configurations C0 (the ordered state), and C1, C2, C3 (thermalized from 3 different random starts). Also shown are results for the KBD algorithm with one Metropolis sweep per iteration (KBD+M). The averages were taken over at least 50,000 iterations, and the statistical error in the last digit is given in brackets. For the two smallest lattice sizes we also computed the exact result.

L=2  L=4  L=8  L=16  L=32  L=64
C0  0.4999(5) 0.3751(3) 0.2446(3) 0.1055(6) 0.1061(10) 0.0718(8)
C1  0.3751(3) 0.2811(3) 0.4532(3) 0.1576(11) 0.0946(16) 0.0768(12)
C2  —  0.4532(3) 0.0977(1) 0.1545(16) 0.1036(17) 0.0714(9)
C3  —  0.5625(0) 0.2299(2) 0.1332(11) 0.1107(10) 0.0771(8)
KBD+M 0.3751(4) 0.2720(3) 0.2020(2) 0.1454(2)  0.1033(2)  0.0731(1)
exact  0.3750000 0.2720588  —  —  —  —